One person can do a job in 8 hours. a second person can do it in 12 hours. if the first peron works 2 hours less than the second, how many hours will it take them working together?
Can you please show me how to do the work I don't need the answer persay just help with setting up the problem! This would be so appreciated. 

Answer by scott8148(6628)

the second person works 2 hr MORE and completes 1/6 (2/12) of the job on their own
___ leaving 5/6 of the job to complete together

let x= time working together ___ they each complete some fraction of the job based on their individual times

(x / 8) + (x / 12) = 5/6


this could also be written as a single step ___ (x / 8) + [(x+2) / 12] = 1